OpenStudy (anonymous):

need help on this one: A leather coin purse has the shape of a triangle. Two sides are equal in length and the third side is 3 centimeters shorter than one and one half times the length of the equal sides. The perimeter is 28.5 centimeters. Find the lengths of the sides.

OpenStudy (anonymous):

let x be equal sides and y be remaining side given the info, y=1.5x-3 and the perimeter of the triangle is x+x+y=28.5 =x+x+1.5x-3=28.5 =x+x+1.5x=31.5 =3.5x=31.5 =x/3.5=31.3/3.5 x=9 y= 1.5 X 9-3=10.5